### v5

# Study on data analysis for Ives-Stilwell-type experiments

Abstract
Ives-Stilwell experiment in 1938 was a historic experiment for confirming Einstein’s special relativity, and various modern types have been repeated by use of laser technology. In this paper, we reveal and solve a long-lasting unrecognized fundamental issue that the data analysis for all those experiments was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. Based on existing data, we propose a justified data analysis and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus resulting in a great advance in the experimental verification of Lorentz invariance via the Doppler effect.
pacs:
03.30.+p, 03.50.De
I Introduction
The relativistic Doppler effect is one of the basic results of Einstein’s special relativity, and it states that a plane light wave or a single laser beam or a single photon in free space, observed in any inertial frames of reference, propagates at the same speed but it may exhibit different frequencies and aberrations r1 . The 1938 Ives-Stilwell experiment r2 was a historic research work for confirming Einstein’s special relativity via the Doppler effect, and various modern types of the experiment r3 ; r4 ; r5 ; r6 ; r7 have been repeated by use of laser technology. The concluding result from an Ives-Stilwell-type experiment is presented in the Letter r7 by an international group of collaborators after 15 years of efforts r8 . However the data analysis method for all those experiments r2 ; r3 ; r4 ; r5 ; r6 ; r7 was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. Here, based on existing experimental data r7 , we propose a justified analysis method and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus successfully revealing and solving a long-lasting unrecognized fundamental issue in the experimental confirmation of special relativity.
II Definitions of measurement accuracies and analysis of experimental data
In the excellent Letter, Botermann et al. r7 report that the measurement accuracy of relativistic Doppler effect has reached the order of $10^{-9}$ in an Ives-Stilwell-type experiment, where a moving high-energy ion beam is set at resonance with two laser beams aligned parallel and antiparallel relative to the ion beam, and the ions with two very close transition frequencies behave as Doppler-shifted laser frequency detector.
However we find that, according to a justified data analysis proposed here, the accuracy obtained by the authors r7 is not the accuracy for a single laser beam. The accuracy for a single laser beam is much lower, in the order of $10^{-4}$, different from the accuracy reported in r7 by five orders of magnitude. The analysis is given below.
In their Letter, the definition of measurement accuracy and reported result are given by r7
$\varepsilon$ $=\sqrt{\frac{\nu_{a}\nu_{p}}{\nu_{1}\nu_{2}}}-1,$ (definition) (1)
$\varepsilon$ $=(1.5\pm 2.3)\times 10^{-9},$ (reported  result) (2)
where $\nu_{a}$ and $\nu_{p}$ are, respectively, the frequencies of the two laser beams; $\nu_{1}$ and $\nu_{2}$ are the transition frequencies of ions; $(\nu_{a}\nu_{p}/\nu_{1}\nu_{2})^{1/2}$ is the measured value, with its true value equal to unity. Mathematically, this definition itself is impeccable, which refers to how a measured value is close to its true value: $accuracy=(x_{\textrm{measured}}-x_{\textrm{true}})/x_{\textrm{true}}$, where $x$ is a physical quantity to be examined. However physically, what Eq. (1) describes is an expression of the coupling of various accuracies for two individual laser beams, and it cannot represent the accuracy for any single beam of them. Thus Eq. (1) is not consistent with Einstein’s definition of the Doppler effect that refers to a single laser beam or a single photon which, observed in any inertial frames in free space, propagates at the same speed but may exhibit different frequencies and aberrations r1 . If the measurement accuracy for a single laser beam cannot be obtained, then it is not persuasive and convincing to declare that the Lorentz invariance or special relativity has been verified via the relativistic Doppler effect.
In the Ives-Stilwell-type experiment r7 , observed in the ion-rest frame there are two measured frequencies (namely transition frequencies): $\nu_{1}=546,455,143.0$ MHz and $\nu_{2}=546,474,960.7$ MHz. Because the two frequencies are too close plus inevitable systematic errors, it is impossible to identify which transition frequency responds to which laser beam. Thus we have to evaluate the accuracies respectively for the two measured frequencies for each laser beam.
According to Einstein r1 , for any measured laser beam frequencies $\nu_{a}$ and $\nu_{p}$, and measured ion’s velocity $\beta c$ (with $c$ the vacuum light speed), we have Doppler-shifted laser frequencies observed in the ion-rest frame, given by
$\nu_{a\textrm{-sh}}$ $=\gamma(1+\beta)\nu_{a},$ (antiparallel beam) (3)
$\nu_{p\textrm{-sh}}$ $=\gamma(1-\beta)\nu_{p},$ (parallel beam) (4)
where $\gamma=(1-\beta^{2})^{-1/2}$ is the relativistic factor. Note that either $\nu_{a\textrm{-sh}}$ or $\nu_{p\textrm{-sh}}$ can be taken as the true value of measured $\nu_{1}$ or $\nu_{2}$.
Thus we have two accuracies for measured frequency $\nu_{1}$, given by
$\varepsilon_{a1}$ $=\frac{\nu_{1}-\nu_{a\textrm{-sh}}}{\nu_{a\textrm{-sh}}},$ (antiparallel beam) (5)
$\varepsilon_{p1}$ $=\frac{\nu_{1}-\nu_{p\textrm{-sh}}}{\nu_{p\textrm{-sh}}},$ (parallel beam) (6)
and two accuracies for measured frequency $\nu_{2}$, given by
$\varepsilon_{a2}$ $=\frac{\nu_{2}-\nu_{a\textrm{-sh}}}{\nu_{a\textrm{-sh}}},$ (antiparallel beam) (7)
$\varepsilon_{p2}$ $=\frac{\nu_{2}-\nu_{p\textrm{-sh}}}{\nu_{p\textrm{-sh}}}.$ (parallel beam) (8)
From Eqs. (3) and (4), we have $\nu_{a}\nu_{p}=\nu_{a\textrm{-sh}}\nu_{p\textrm{-sh}}$. Thus Eq. (1) can be written as
$\varepsilon=\sqrt{\frac{\nu_{a\textrm{-sh}}\nu_{p\textrm{-sh}}}{\nu_{1}\nu_{2}}}-1.$ (9)
From Eq. (5) we have $\nu_{1}/\nu_{a\textrm{-sh}}=1+\varepsilon_{a1}$ and from Eq. (8) we have $\nu_{2}/\nu_{p\textrm{-sh}}=1+\varepsilon_{p2}$. Inserting them into Eq. (9) we have
$\varepsilon=\frac{1}{\sqrt{\big{(}1+\varepsilon_{a1}\big{)}\big{(}1+\varepsilon_{p2}\big{)}}}-1.$ (10)
From Eq. (6) we have $\nu_{1}/\nu_{p\textrm{-sh}}=1+\varepsilon_{p1}$ and from Eq. (7) we have $\nu_{2}/\nu_{a\textrm{-sh}}=1+\varepsilon_{a2}$. Inserting them into Eq. (9) we have
$\varepsilon=\frac{1}{\sqrt{\big{(}1+\varepsilon_{p1}\big{)}\big{(}1+\varepsilon_{a2}\big{)}}}-1.$ (11)
According to the Letter r7 , the measured data are given by $\beta=0.338$, $\nu_{1}=546,455,143.0$ MHz, $\nu_{2}=546,474,960.7$ MHz, $\nu_{a}=384,225,534.98$ MHz ($\lambda_{a}=780$ nm), and $\nu_{p}=777,210,326.98$ MHz ($\lambda_{p}=386$ nm). Using these data, from Eqs. (3)-(8) we obtain the accuracies for individual laser beams, given by
$\varepsilon_{a1}$ $=+3.897,464,375,146\times 10^{-4},$ (12)
$\varepsilon_{p1}$ $=-4.258,480,215,884\times 10^{-4},$ (13)
$\varepsilon_{a2}$ $=+4.260,264,947,745\times 10^{-4},$ (14)
$\varepsilon_{p2}$ $=-3.895,975,426,142\times 10^{-4}.$ (15)
Thus from Eqs. (10) and (11) we have
$\varepsilon$ $\approx-\frac{1}{2}(\varepsilon_{a1}+\varepsilon_{p2}+\varepsilon_{a1}\varepsilon_{p2})=1.47\times 10^{-9},$ (16)
$\varepsilon$ $\approx-\frac{1}{2}(\varepsilon_{p1}+\varepsilon_{a2}+\varepsilon_{p1}\varepsilon_{a2})=1.47\times 10^{-9}.$ (17)
From Eqs. (12)-(15) we can see that the measurement accuracies of Doppler effect for individual laser beams are in the order of $10^{-4}$, much lower than the accuracy of about $10^{-9}$ reported in the Letter r7 .
From Eq. (16) we find that the accuracy $\varepsilon$, defined in r7 , is a combination of $\varepsilon_{a1}$ (the accuracy for antiparallel laser beam for $\nu_{1}$) and $\varepsilon_{p2}$ (the accuracy for parallel laser beam for $\nu_{2}$). $\varepsilon_{a1}$ and $\varepsilon_{p2}$ almost cancel each other to become in the same order of magnitude as $\varepsilon_{a1}\varepsilon_{p2}$ with opposite signs, namely $\varepsilon_{a1}+\varepsilon_{p2}=148.8949\times 10^{-9}$ and $\varepsilon_{a1}\varepsilon_{p2}=-151.8443\times 10^{-9}$, so that $\varepsilon=1.47\times 10^{-9}$ is obtained. The same thing happens to Eq. (17).
In fact, it is easy to understand mathematically that the accuracies for single laser beams are essentially different from the coupling accuracy. For example, from Eq. (10) we see that even when $\varepsilon=0$ holds, we can have $\varepsilon_{a1}=-0.5\,(1)$ and $\varepsilon_{p2}=1\,(-0.5)$ holding, not necessarily $\varepsilon_{a1}=0$ and $\varepsilon_{p2}=0$.
III Conclusions and remarks
In this paper, we have revealed and solved a long-lasting unrecognized fundamental issue on data analysis for Ives-Stilwell-type experiments. We have shown that $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$, given by Eq. (1), is an expression of the coupling of various accuracies for two individual laser beams, and it cannot represent the accuracy for any single beam of them, namely it cannot represent the measurement accuracy of the Doppler effect defined by Einstein. To confirm this, by use of the experimental data provided in the Letter where the measured ion beam velocity is given by $\beta=0.338$ r7 , we calculated all four accuracies for single laser beams, showing that the accuracy for a single laser beam is much lower, in the order of $10^{-4}$, different from the accuracy reported in the Letter by five orders of magnitude.
Einstein’s definition of the relativistic Doppler effect is widely accepted by the physics community, and it is a scientific consensus, and of course, it is within the scope of the body of knowledge available.
[…]
One might criticize that, the analysis in the paper is based mistakenly on an approximate value of the ion beam’s velocity ($\beta=0.338$), and the actual velocity of the ions contributing to the [relativistic Doppler] resonance signal can be determined to be $\beta=0.338,377,22$. If $\beta=0.338,377,22$ is used in the paper, the supposed discrepancy disappears.
However, if the actual velocity of the ions can be determined by applying the relativistic Doppler formula to the Ives-Stilwell-type experiment, which is designed for confirming the relativistic Doppler formula itself, then the Ives-Stilwell-type experiment must be required to support both the confirmation and application of the Doppler formula at the same time. This is apparently a Circular Reasoning or Logical Fallacy. That is because: (i) to confirm the Doppler formula, it is necessary to assume that the ion velocity $\beta$ has been measured; (ii) conversely, to measure the ion velocity $\beta$, it is necessary to assume that the Doppler formula has been confirmed. Thus this criticism is not appropriate.
One might argue that the error reported in the paper simply indicates that $\beta~{}[=0.338,377,22]$ is slightly different from the value $[\beta=0.338]$ given in the Letter r7 . However this is not true. The error is that the quantity $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ cannot represent the measurement accuracy of Einstein’s Doppler effect, because Einstein’s Doppler effect refers to the effect of a single laser beam.
It should be noted that, although $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ is not dependent on $\beta$, the measurement accuracy for a single laser beam is sensitively dependent on $\beta$; for example, $\varepsilon_{a2}=4.260,264,947,745\times 10^{-4}$ for $\beta=0.338$, while $\varepsilon_{a2}=9.432,048,656,75\times 10^{-10}$ for $\beta=0.338,377,22$, with the measurement accuracy changed by five (5) orders of magnitude for a slight $\beta$-change of about $0.1\%$.
One might argue that although the measurement accuracy of the Doppler effect can only be represented by that from a single laser beam, it is possible to correlate the two measurements from antiparallel and parallel laser beams in order to improve accuracy without challenging the physical phenomenon. However this is not true, because the Doppler phenomenon is judged by checking the accuracy of Doppler effect from a single laser beam, and the correlation of the two measurements may change the accuracy, and further challenge the judgment of Doppler phenomenon.
Generally speaking, the experiment itself does not prove anything, as it only provides readings of instruments and meters. It is the researchers of the experiment who use these readings to form a chain of evidence to support their conjectures. Therefore, for the same readings of instruments and meters, the formed chain of evidence may depend on the way of thinking of the researchers, resulting in different conclusions. The Ives-Stilwell experiment is probably a typical example, as shown below.
In 1938, Ives and Stilwell famously reported that the results of their experiment “appear to be a satisfactory confirmation of the Larmor-Lorentz theory” r2 , namely their experiment satisfactorily confirmed Lorentz’s ether theory. However Jones argued that the predictions obtained from Einstein’s special relativity are identical with those obtained by Ives based on Lorentz’s ether theory r9 , namely the Ives-Stilwell experiment satisfactorily confirmed Einstein’s special relativity.
From above it can be seen that Jones’ and Ives-Stilwell chains of evidence, formed based on the same experimental data, support completely different conclusions.
The present paper might be another interesting example, with the same experimental data used to form an alternative chain of evidence, leading to a unique and definite conclusion.
Finally, it should be emphasized that the data analysis method in various Ives-Stilwell-type experiments r3 ; r4 ; r5 ; r6 ; r7 is essentially the same as that used by the first Ives-Stilwell experiment in 1938 r2 , where the measurement accuracies for single light beams were not provided, so in fact, the relativistic Doppler effect was not truly confirmed. Therefore, the present work is the first correct confirmation of the relativistic Doppler effect in the Ives-Stilwell-type experiment based on the data provided by Botermann et al. r7 .
The principle of relativity requires Maxwell equations to have Lorentz invariance, resulting in invariance of the speed of light in free space r10 and relativistic Doppler effect r1 . Thus the present work is a great advance in the experimental verification of Lorentz invariance via the Doppler effect, just like the verification via the invariance of light speed in the astronomical observations of high-energy gamma rays r11 ; r12 .
Disclosures. The author declares no conflicts of interest.
Data availability. Data underlying the results presented in this paper are available within the paper and Supplement 1.
Supplemental document. See Supplement 1 for supporting content r13 .